OpenStudy (anonymous):

State the domain of the rational function. f(x)=13/2-x Please explain the steps to me, if possible!

jimthompson5910 (jim_thompson5910):

if 2-x is the denominator then the denominator cannot be 0 if it was 0, then 2-x = 0 2-x+x = 0+x 2+0x = x 2 = x x = 2 which means that x = 2 makes 2-x equal to zero so we must kick out the number 2 from the domain to avoid division by zero

jimthompson5910 (jim_thompson5910):

the domain is therefore: the set of all real numbers but x cannot equal 2

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Just positive 2, no -2?

jimthompson5910 (jim_thompson5910):

positive 2

jimthompson5910 (jim_thompson5910):

-2 works because it does not make the denominator 0

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

sure thing